tom is 5 years old, his dad is 5 x toms age. in how many years time will tom be half his fathers age
Answer by graphmatics(170)

We are given that tom is 5. As his dad is 5 times tom's age his dad is 25. let t be the number of years we seek then in these t years tom will be 5+t years old and his dad will be 25+t years old. We have
(5+t)/(25+t)=1/2
(5+t)*2=25+t
10+2*t=25+t
2*t-t=25-10
t=15